# Armijo Line-search Makes (Stochastic) Gradient Descent Go Fast

### v1

@@ 7 Conclusion
We explored the theoretical properties of GD-LS for a class of functions satisfying non-uniform smoothness. For a range of practically useful convex and non-convex functions, we proved that Armijo-LS can effectively adapt to the objective’s structural properties and enable faster convergence for GD. In particular, we showed that GD-LS can either match or provably improve upon the sublinear rate of GD(1/L), and do so without relying on the knowledge of problem-dependent constants. Furthermore, for specific problems in supervised learning and reinforcement learning, we demonstrated that GD-LS can match the fast convergence of algorithms tailored for these problems. In conclusion, our results show the universality and effectiveness of GD-LS. We believe that investigating the behaviour of GD-LS for a broader class of non-convex functions, and formally characterizing the advantage of Armijo-LS for other algorithms (such as Nesterov accelerated gradient) remain important directions for future work.
@@ Acknowledgments
We thank Damien Scieur, Mark Schmidt, Curtis Fox, Michael Lu and Siyi Meng for helpful discussions and feedback, and Anh Dang for help with the initial experiments.
@@
Appendix A Proofs for Section 2
In order to prove that commonly used functions in machine learning satisfy the assumptions in Section 2, we will require some additional assumptions and intermediate results.
@@ Appendix A Proofs for Section 2
$f$ is twice-differentiable and satisfies the following non-uniform smoothness property: for constants $L_{c}$ , $L_{g}>0$,
$\left\|\nabla^{2}f(\theta)\right\|\leq L_{c}+L_{g}\,\left\|\nabla f(\theta)\right\|$
Unlike Assumption 2Assumption 5 corresponds to the standard non-uniform smoothness assumption made in the literature (Zhang et al., 2019, 2020) and implies the following result.
Lemma 4 (Lemma A3 (Zhang et al., 2020)).
If $f$ satisfies Assumption 5, then the following inequality holds for all $x,y$ such that $\left\|x-y\right\|\leq\frac{q}{L_{g}}$ where $q>0$ is a constant,
@@
where $A:=1+e^{q}-\frac{e^{q}-1}{q}$ and $B:=\frac{e^{q}-1}{q}$.
@@
For a finite-sum objective, $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$, if, for all $i$, $f_{i}$ satisfies Assumption 3 with the constants equal to $\nu,\omega$, then,
$\left\|\nabla f(\theta)\right\|\leq\nu\,f(\theta)+\omega$
Furthermore, if for all $i$, $f_{i}$ also satisfies Assumption 5 with constants $L_{c},L_{g}$, then, the following inequalities hold $L_{0}:=L_{c}+L_{g}\,\omega$ and $L_{1}:=\nu\,L_{g}$:
(a) for all $\theta$,
$\left\|\nabla^{2}f(\theta)\right\|\leq L_{0}+L_{1}\,f(\theta)$
(b) for all $x,y$ such that $\left\|x-y\right\|\leq\frac{q}{L_{1}}$ where $q\geq 1$ is a constant,
$f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{AL_{0}+B\,L_{1}\,f(x)}{2}\,\left\|y-x\right\|_{2}^{2}\,,$
where $A:=1+e^{q}-\frac{e^{q}-1}{q}$ and $B:=\frac{e^{q}-1}{q}$.
Hence, if $f_{i}(\theta)$ satisfies Assumption 3 and Assumption 5, then $f(\theta)$ satisfies Assumption 2 and Assumption 3.
@@
For the first part, note that,
@@
For the second part,
$\left\|\nabla^{2}f(\theta)\right\|$ $=\left\|\frac{1}{n}\,\sum_{i}\nabla^{2}f_{i}(\theta)\right\|\leq\frac{1}{n}\,\sum_{i}\left\|\nabla^{2}f_{i}(\theta)\right\|\leq L_{c}+\frac{L_{g}}{n}\,\sum_{i}\left\|\nabla f_{i}(\theta)\right\|$
$\leq L_{c}+\frac{L_{g}}{n}\,\sum_{i}[\nu\,f_{i}(\theta)+\omega]=[L_{c}+L_{g}\,\omega]+\nu\,L_{g}\,f(\theta)$
For the third part, using Lemma 5, for all $i$,
$f_{i}(y)$ $\leq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{AL_{c}+B\,L_{g}\,\left\|\nabla f_{i}(x)\right\|}{2}\left\|y-x\right\|_{2}^{2}$
$\leq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{AL_{c}+B\,L_{g}\,[\nu\,f_{i}(x)+\omega]}{2}\left\|y-x\right\|_{2}^{2}$ (Since $f_{i}$ satisfies Assumption 3 with constants $\nu,\omega$)
$=f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{A\left(L_{c}+\frac{B}{A}\,L_{g}\,\omega\right)+B\,L_{g}\,\nu\,f_{i}(x)}{2}\left\|y-x\right\|_{2}^{2}$
$\leq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{A\left(L_{c}+L_{g}\,\omega\right)+B\,L_{g}\,\nu\,f_{i}(x)}{2}\left\|y-x\right\|_{2}^{2}$ (Since $B\leq A$)
Summing the LHS and RHS for $i=1$ to $n$ and dividing by $n$ completes the proof. ∎
Hence, if the conditions of Lemma 5 are satisfied, then the non-uniform smoothness condition in Assumption 2 is satisfied. The following propositions show that such a non-uniform smoothness condition is satisfied for linear logistic regression, linear model with an exponential loss, linear multi-class classification using the cross-entropy loss, generalized linear models with a logistic link function and the softmax policy gradient objective for multi-armed bandits and tabular MDPs.
@@
Clearly, $f_{i}(\theta)\geq 0$ and hence $f(\theta)\geq 0$ for all $\theta$. $f(\theta)$ is a finite-sum objective. Calculating the gradient and hessian for $f_{i}(\theta):=\ln(1+\exp(-y_{i}\langle x_{i},\theta\rangle)$,
@@
$\left\|\nabla f_{i}(\theta)\right\|$ $=\frac{\exp(-y_{i}\langle x_{i},\theta\rangle)}{1+\exp(-y_{i}\langle\,x_{i},\theta\rangle)}\left\|x_{i}\right\|$
$\implies\left\|\nabla f_{i}(\theta)\right\|$ $\leq\left\|x_{i}\right\|\,\ln\left(1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)\right)=\left\|x_{i}\right\|\,f_{i}(\theta)$ (For all $x$, $\frac{x}{1+x}\leq\ln(1+x)$)
Hence, for all $i$, $f_{i}$ satisfies Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. Bounding the Hessian,
@@
Hence, for all $i$, $f_{i}$ satisfies Assumption 5 with $L_{c}=0$ and $L_{g}=\max_{i}\left\|x_{i}\right\|$. Using Lemma 5, we conclude that $f(\theta)$ satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=L_{g}\,\nu=\max_{i}\left\|x_{i}\right\|_{2}^{2}$, and Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. ∎
@@
$\implies\left\|\nabla f_{i}(\theta)\right\|$ $=|\pi_{i}(\theta)-y_{i}|\,\pi_{i}(\theta)\,\underbrace{(1-\pi_{i}(\theta))}_{\leq 1}\,\left\|x_{i}\right\|\leq|\pi_{i}(\theta)-y_{i}|\,|\pi_{i}(\theta)-y_{i}+y_{i}|\,\left\|x_{i}\right\|$
$\leq\left[|\pi_{i}(\theta)-y_{i}|\,|\pi_{i}(\theta)-y_{i}|+y_{i}\,|\pi_{i}(\theta)-y_{i}|\right]\,\left\|x_{i}\right\|$ (Triangle inequality)
$=2\,\left\|x_{i}\right\|\,\left[\frac{1}{2}\,\left(\pi_{i}(\theta)-y_{i}\right)^{2}\right]+\left\|x_{i}\right\|$ (Since $y_{i}\,|\pi_{i}(\theta)-y_{i}|\in[0,1]$)
$\implies\left\|\nabla f_{i}(\theta)\right\|$ $\leq 2\,\left\|x_{i}\right\|\,f_{i}(\theta)+\left\|x_{i}\right\|$
Hence, for all $i$, $f_{i}$ satisfies Assumption 3 with $\nu=2\,\max_{i\in[n]}\left\|x_{i}\right\|$ and $\omega=\max_{i\in[n]}\left\|x_{i}\right\|$. Calculating the Hessian,
@@
Hence, for all $i$, $f_{i}(\theta)$ satisfies Assumption 5 with $L_{c}=\frac{1}{16}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and $L_{g}=\max_{i\in[n]}\left\|x_{i}\right\|$. Using Lemma 5, we conclude that $f(\theta)$ satisfies Assumption 2 with $L_{0}=\frac{17}{16}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and $L_{1}=2\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, and Assumption 3 with $\nu=2\max_{i}\left\|x_{i}\right\|$ and $\omega=\max_{i}\left\|x_{i}\right\|$. ∎
@@
where $\theta_{i}\in\mathbb{R}^{d}$ for $i\in[C]$ and $\theta=[\theta_{1},\theta_{2},\ldots,\theta_{C}]$. Multi-class logistic regression satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=4\,\max_{m\in[n]}\left\|x\right\|^{2}_{1}$, and Assumption 3 with $\nu=2\max_{i}\left\|x_{i}\right\|$ and $\omega=0$.
@@
Let us consider a single input-output pair $(x,y)$ and calculate the gradient for a single function in the finite-sum. Define $\ell(\theta):=\text{KL}(y||\pi_{\theta})$ where $y$ is a $C$-dimensional one-hot vector, $x\in\mathbb{R}^{d}$ and $\pi_{\theta}\in\Delta_{C}$ s.t. $\pi_{\theta}(i)=\frac{\exp(\langle x,\theta_{i}\rangle)}{\sum_{k=1}^{C}\exp(\langle x,\theta_{k}\rangle)}$
$\frac{\partial\ell(\theta)}{\partial\theta_{i}}$ $=[\pi_{\theta}(i)-y_{i}]\,x\implies\left\|\frac{\partial\ell(\theta)}{\partial\theta_{i}}\right\|_{1}=|\pi_{\theta}(i)-y_{i}|\,\left\|x\right\|_{1}$
$\implies\left\|\nabla_{\theta}\ell(\theta)\right\|_{1}$ $=\left\|x\right\|_{1}\,\sum_{i=1}^{C}|\pi_{\theta}(i)-y_{i}|$
Since $y$ is a one-hot vector, let $i^{*}$ be the index corresponding to the non-zero entry. Hence, $y_{i^{*}}=1$ and for all $j\neq i^{*}$, $y_{j}=0$. With this,
$\leq\left\|x\right\|_{1}\,\sum_{i\neq i^{*}}\pi_{\theta}(i)+[1-\pi_{\theta}(i^{*})]=2\,[1-\pi_{\theta}(i^{*})]\,\left\|x\right\|_{1}$
$\leq 2\,\left\|x\right\|_{1}\,\ln\left(\frac{1}{\pi_{\theta}(i^{*})}\right)$ (For all $z\in[0,1]$, $1-z\leq\ln(1/z)$)
$=2\,\left\|x\right\|_{1}\,\sum_{i=1}^{C}y_{i}\,\ln\left(\frac{y_{i}}{\pi_{\theta}(i)}\right)$ (Using that $y$ is a one-hot vector)
$\implies\left\|\nabla_{\theta}\ell(\theta)\right\|_{1}$ $\leq 2\,\left\|x\right\|_{1}\,\text{KL}(y||\pi_{\theta})=2\,\left\|x\right\|_{1}\,\ell(\theta)\,.$
Hence, $\ell(\theta)$ satisfies Assumption 3 with $\nu=2\left\|x\right\|_{1}$ and $\omega=0$. Let us now bound the Hessian. The Hessian can be written as a Kronecker product of a $C\times C$ matrix which corresponds to the Jacobian of the softmax function, and a $d\times d$ rank-one matrix formed using the features. Specifically,
@@
Hence, for a single $(x,y)$ pair, we can conclude that, $\ell(\theta)$ satisfies Assumption 5 with $L_{g}=2\left\|x\right\|_{1}$.
Combining the above results implies that for all $m\in[n]$, $\text{KL}(y^{m}||\pi^{m}_{\theta})$ satisfies  Assumption 3 with $\nu=2\max_{m\in[n]}\left\|x\right\|_{1}$ and Assumption 5 with $L_{g}=2\max_{m\in[n]}\left\|x\right\|_{1}$. Since $f$ is a finite-sum, we use Lemma 5 to conclude that $f(\theta)$ satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=4\,\max_{m\in[n]}\left\|x\right\|^{2}_{1}$, and Assumption 3 with $\nu=2\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. ∎
@@
satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, and Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$.
@@
Clearly, $f_{i}(\theta)\geq 0$ and hence $f(\theta)\geq 0$ for all $\theta$. $f(\theta)$ is a finite-sum objective. Calculating the gradient and hessian for $f_{i}(\theta):=\exp(-y_{i}\langle x_{i},\theta\rangle)$,
$\nabla f_{i}(\theta)$ $=-\exp(-y_{i}\langle x_{i},\theta\rangle)y_{i}\,x_{i}\quad\text{;}\quad\nabla^{2}f_{i}(\theta)=\exp(-y_{i}\langle x_{i},\theta\rangle)\,y_{i}^{2}\,x_{i}\,x_{i}^{T}$
$\implies\left\|\nabla f_{i}(\theta)\right\|$ $=\exp(-y_{i}\langle x_{i},\theta\rangle)\,\left\|x_{i}\right\|=f_{i}(\theta)\,\left\|x_{i}\right\|$
Hence, for all $i$, $f_{i}$ satisfies Assumption 3 with $\nu=\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. Bounding the Hessian,
@@
Hence, for all $i$, $f_{i}$ satisfies Assumption 5 with $L_{0}=0$ and $L_{g}=\max_{i}\left\|x_{i}\right\|$. Since $f$ is a finite-sum objective, using Lemma 5, we conclude that $f(\theta)$ satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=L_{g}\,\nu=\max_{i}\left\|x_{i}\right\|_{2}^{2}$. ∎
See 3
@@
From Mei et al. (2020, Lemma 17), we know that, if $\Delta:=\min_{a\neq a^{*}}r(a^{*})-r(a)$ is the minimum reward gap, then,
$\left\|\nabla\ell(\theta)\right\|=\left\|\nabla_{\theta}\langle\pi_{\theta},r\rangle\right\|$ $\leq\frac{\sqrt{2}}{\Delta}\,\ell(\theta)$
Hence, the loss for the bandit problem satisfies Assumption 3 with $\nu=\frac{\sqrt{2}}{\Delta}$ and $\omega=0$. From Mei et al. (2021, Lemma 2), we know that
$\left\|\nabla^{2}\ell(\theta)\right\|$ $=\left\|\nabla^{2}\langle\pi_{\theta},r\rangle\right\|\leq 3\,\left\|\nabla\langle\pi_{\theta},r\rangle\right\|=\left\|\nabla\ell(\theta)\right\|$
Hence, the loss for the bandit problem satisfies Assumption 5 with $L_{c}=0$ and $L_{g}=3$. Using Lemma 5 with $n=1$, we can conclude the the loss for the bandit problem satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\frac{3\sqrt{2}}{\Delta}$. From Mei et al. (2020, Lemma 3), we know that,
@@
Hence, the loss for the bandit problem satisfies Assumption 4 with $\mu(\theta)=\pi_{\theta}(a^{*})$. ∎
@@
where $\pi^{*}$ is the optimal policy. $f(\theta)$ satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\,\frac{\sqrt{2}}{\Delta\,(1-\gamma)}$,  Assumption 3 with $\nu=\frac{\sqrt{2}}{\Delta\,(1-\gamma)}$ and $\omega=0$ and  Assumption 4 with $\mu(\theta)=\frac{\min_{s}\pi_{\theta}(\pi^{*}(s)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}}\rho(s)}$.
@@
From Mei et al. (2020, Lemma 28), we know that if $\Delta:=\min_{s\in{\mathcal{S}}}Q^{*}(s,\pi^{*}(s))-\max_{a\neq\pi^{*}(s)}Q(s,a)$, then,
$\left\|\nabla f(\theta)\right\|\leq\frac{\sqrt{2}}{\Delta\,(1-\gamma)}\,f(\theta)$
Hence, the loss for the tabular MDP problem satisfies Assumption 3 with $\nu=\frac{\sqrt{2}}{\Delta\,(1-\gamma)}$ and $\omega=0$. Assuming that the starting state distribution has full support, i.e. $\rho(s)>0$, from Mei et al. (2020, Lemma 6), we know that,

### v2

@@ 7 Conclusion
We analyzed GD-LS for a class of functions satisfying non-uniform smoothness. For a range of practical convex and non-convex functions, we proved that Armijo-LS can enable GD to adapt to the objective’s properties and result in faster convergence. In particular, we showed that, for specific problems in machine learning, GD-LS can (i) either match or provably improve upon the sublinear rate of GD(1/L), (ii) do so without relying on the knowledge of problem-dependent constants and (iii) match the fast convergence of algorithms tailored for these problems. Our results thus show the universal effectiveness of GD-LS.
We believe that analyzing GD-LS for a broader class of non-convex functions, and characterizing the advantage of using Armijo-LS for other algorithms (such as Nesterov accelerated gradient) are important future directions. We also plan to investigate whether other adaptive step-size schemes such as AdaGrad (Duchi et al., 2011) or Adam (Kingma & Ba, 2014) can also provably adapt to the non-uniform smoothness and result in faster convergence rates.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgments
We thank Damien Scieur, Mark Schmidt, Curtis Fox, Michael Lu and Siyi Meng for helpful discussions and feedback, and Anh Dang for help with the initial experiments. This research was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant RGPIN-2022-04816.
@@
Appendix A Proofs for Sec. 2
@@
If $f$ is twice differentiable and $(L_{c},L_{g})$-non-uniform smooth as defined in Zhang et al. (2019), then, for constants $L_{c}$ , $L_{g}>0$,
$\left\|\nabla^{2}f(\theta)\right\|\leq L_{c}+L_{g}\,\left\|\nabla f(\theta)\right\|.$ (10)
See 3
Proof.
Using Li et al. (2023, Lemma 3.5), we know that $(L_{c},L_{g})$ non-uniform smoothness also implies that,
$\left\|\nabla f(\theta)\right\|_{2}^{2}$ $\leq 2\left[L_{c}+2L_{g}\,\left\|\nabla f(\theta)\right\|\right]\,(f(\theta)-f^{*})$
$\leq 2\left[L_{c}+2L_{g}\,\left\|\nabla f(\theta)\right\|\right]\,f(\theta)$ (Since $f$ is non-negative)
$\implies\left\|\nabla f(\theta)\right\|$ $\leq 4L_{g}\,f(\theta)+\sqrt{2L_{c}}\,\sqrt{f(\theta)}$ (By completing the square w.r.t $\left\|\nabla f(\theta)\right\|$)
Consider two cases depending on whether $f(\theta)$ is larger than 1.
Case 1: If $f(\theta)\leq 1\implies\sqrt{f(\theta)}\leq 1$. The above inequality can be simplified as: $\left\|\nabla f(\theta)\right\|\leq 4L_{g}\,f(\theta)+\sqrt{2L_{c}}$.
Case 2: If $f(\theta)>1\implies\sqrt{f(\theta)}\leq f(\theta)$. The above inequality can be simplified as: $\left\|\nabla f(\theta)\right\|\leq 4L_{g}\,f(\theta)+\sqrt{2L_{c}}\,f(\theta)$.
Combining both cases, we have that,
$\left\|\nabla f(\theta)\right\|$ $\leq\left(8L_{g}+\sqrt{2L_{c}}\right)f(\theta)+\sqrt{2L_{c}}$ (11)
Hence, Assn. 3 is satisfied with $\nu=8L_{g}+\sqrt{2L_{c}}$ and $\omega=\sqrt{2L_{c}}$. Using this result in the definition of non-uniform smoothness in Assn. 5,
$\left\|\nabla^{2}f(\theta)\right\|$ $\leq L_{c}+L_{g}\,\left[\left(8L_{g}+\sqrt{2L_{c}}\right)f(\theta)+\sqrt{2L_{c}}\right]=\left(L_{c}+L_{g}\,\sqrt{2L_{c}}\right)+L_{g}\left(8L_{g}+\sqrt{2L_{c}}\right)f(\theta)$ (12)
Hence, Assn. 2 (b) is satisfied with $L_{0}=L_{c}+L_{g}\,\sqrt{2L_{c}}$ and $L_{1}=L_{g}\left(8L_{g}+\sqrt{2L_{c}}\right)$.
Using Zhang et al. (2020, Lemma A.3), if $f$ satisfies Assn. 5, then, for all $x,y$ such that $\left\|x-y\right\|\leq\frac{q}{L_{g}}$ where $q>0$ is a constant,
@@
where $A:=1+e^{q}-\frac{e^{q}-1}{q}$ and $B:=\frac{e^{q}-1}{q}$. Combining the above inequality with Eq. 11,
$f(y)$ $\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{AL_{c}+B\,L_{g}\,[\left(8L_{g}+\sqrt{2L_{c}}\right)f(x)+\sqrt{2L_{c}}]}{2}\left\|y-x\right\|_{2}^{2}$
$=f(x)+\langle\nabla f(x),y-x\rangle+\frac{A\left(L_{c}+\frac{B}{A}\,L_{g}\,\sqrt{2L_{c}}\right)+B\,L_{g}\,\left(8L_{g}+\sqrt{2L_{c}}\right)\,f(x)}{2}\left\|y-x\right\|_{2}^{2}$
$\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{A\left(L_{c}+L_{g}\,\sqrt{2L_{c}}\right)+B\,L_{g}\,\left(8L_{g}+\sqrt{2L_{c}}\right)\,f(x)}{2}\left\|y-x\right\|_{2}^{2}$ (Since $B\leq A$)
Hence, Assn. 2 (a) is satisfied with $L_{0}=L_{c}+L_{g}\,\sqrt{2L_{c}}$ and $L_{1}=L_{g}\left(8L_{g}+\sqrt{2L_{c}}\right)$. ∎
In order to prove that commonly used functions in machine learning satisfy the assumptions in Sec. 2, we will require the following lemma.
@@
For a finite-sum objective, $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$, if, for all $i$, $f_{i}$ is non-negative and satisfies Assn. 5 with constants $L_{c}$ and $L_{g}$, then, $f$ satisfies Assn. 1, 2 and 3 with constants $L_{0}=L_{c}+L_{g}\,\sqrt{2L_{c}}$, $L_{1}=L_{g}\,\left(8L_{g}+\sqrt{2L_{c}}\right)$, $\nu=8L_{g}+\sqrt{2L_{c}}$ and $\omega=\sqrt{2L_{c}}$.
@@
If $f_{i}$ is non-negative for all $i$, then, $f$ is non-negative, thus satisfying Assn. 1. We will first prove that if $f_{i}$ is non-negative and satisfies Assn. 2 and 3, then $f$ also satisfies these assumptions with the same constants.
If $f_{i}$ satisfies Assn. 2 (a), then,
$f_{i}(y)$ $\leq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{AL_{0}+B\,L_{1}\,f_{i}(x)}{2}\left\|y-x\right\|_{2}^{2}$
Summing the LHS and RHS for $i=1$ to $n$ and dividing by $n$ completes the proof that $f$ satisfies Assn. 2 (a). If $f_{i}$ satisfies Assn. 2 (b), then,
$\left\|\nabla^{2}f(\theta)\right\|$ $=\left\|\frac{1}{n}\,\sum_{i}\nabla^{2}f_{i}(\theta)\right\|\leq\frac{1}{n}\,\sum_{i}\left\|\nabla^{2}f_{i}(\theta)\right\|\leq L_{0}+\frac{L_{1}}{n}\,\sum_{i}f_{i}(\theta)=L_{0}+L_{1}\,f(\theta)$
Hence, $f$ satisfies Assn. 2 with the same constants $L_{0}$ and $L_{1}$. If $f_{i}$ satisfies Assn. 3, then,
@@
Hence, $f$ satisfies Assn. 3 with the same constants $\nu$ and $\omega$. From Prop. 3, if $f_{i}$ satisfies Assn. 5 with constants $L_{c},L_{g}$, then, $f_{i}$ and consequently $f$ satisfies Assn. 2 and 3 with constants $L_{0}=L_{c}+L_{g}\,\sqrt{2L_{c}}$, $L_{1}=L_{g}\,\left(8L_{g}+\sqrt{2L_{c}}\right)$, $\nu=8L_{g}+\sqrt{2L_{c}}$ and $\omega=\sqrt{2L_{c}}$. ∎
A.1 Examples satisfying Assn. 1, 2 and 3
We will use the above lemma to prove that linear logistic regression, exponential loss with a linear model, linear multi-class classification using the cross-entropy loss, generalized linear models with a logistic link function and the softmax policy gradient objective for multi-armed bandits and tabular MDPs satisfy Assn. 1, 2 and 3.
@@
Clearly, $f_{i}(\theta)\geq 0$ for all $\theta$. Calculating the gradient and hessian for $f_{i}(\theta):=\ln(1+\exp(-y_{i}\langle x_{i},\theta\rangle)$,
@@
Bounding the Hessian,
@@
Hence, for all $i$, $f_{i}$ satisfies Assn. 5 with $L_{c}=0$ and $L_{g}=\max_{i}\left\|x_{i}\right\|$. Using Lemma 4, we conclude that $f(\theta)$ satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=8\,\max_{i}\left\|x_{i}\right\|_{2}^{2}$, and Assn. 3 with $\nu=8\,\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. ∎
@@
Hence, for all $i$, $f_{i}(\theta)$ satisfies Assn. 5 with $L_{c}=\frac{1}{16}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and $L_{g}=\max_{i\in[n]}\left\|x_{i}\right\|$.
Using Lemma 4, we conclude that $f(\theta)$ satisfies Assn. 2 with $L_{0}=\frac{8+\sqrt{2}}{16\sqrt{2}}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$ and $L_{1}=\frac{16\sqrt{2}+1}{\sqrt{8}}\,\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, and Assn. 3 with $\nu=\frac{16\sqrt{2}+1}{\sqrt{8}}\max_{i}\left\|x_{i}\right\|$ and $\omega=\frac{1}{\sqrt{8}}\max_{i}\left\|x_{i}\right\|$. ∎
@@
where $\theta_{i}\in\mathbb{R}^{d}$ for $i\in[C]$ and $\theta=[\theta_{1},\theta_{2},\ldots,\theta_{C}]$. Multi-class logistic regression satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=32\,\max_{m\in[n]}\left\|x\right\|^{2}_{1}$, and Assn. 3 with $\nu=16\max_{i}\left\|x_{i}\right\|$ and $\omega=0$.
@@
Let us consider a single input-output pair $(x,y)$ and calculate the gradient for a single function in the finite-sum. Define $\ell(\theta):=\text{KL}(y||\pi_{\theta})$ where $y$ is a $C$-dimensional one-hot vector, $x\in\mathbb{R}^{d}$ and $\pi_{\theta}\in\Delta_{C}$ s.t. $\pi_{\theta}(i)=\frac{\exp(\langle x,\theta_{i}\rangle)}{\sum_{k=1}^{C}\exp(\langle x,\theta_{k}\rangle)}$. Clearly, $\ell(\theta)\geq 0$. Calculating its gradient and Hessian,
$\frac{\partial\ell(\theta)}{\partial\theta_{i}}$ $=[\pi_{\theta}(i)-y_{i}]\,x\,.$
The Hessian can be written as a Kronecker product of a $C\times C$ matrix which corresponds to the Jacobian of the softmax function, and a $d\times d$ rank-one matrix formed using the features. Specifically,
@@
Hence, for a single $(x,y)$ pair, we can conclude that, $\ell(\theta)$ satisfies Assn. 5 with $L_{g}=2\left\|x\right\|_{1}$. Hence, $f_{i}$ satisfies Assn. 5 with $L_{c}=0$ and $L_{g}=2\,\max_{i\in[n]}\left\|x_{i}\right\|_{1}$.
Using Lemma 4, we can conclude that $f(\theta)$ satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=32\,\max_{i\in[n]}\left\|x_{i}\right\|^{2}_{1}$, and Assn. 3 with $\nu=16\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. ∎
@@
satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=8\max_{i\in[n]}\left\|x_{i}\right\|_{2}^{2}$, and Assn. 3 with $\nu=8\max_{i}\left\|x_{i}\right\|$ and $\omega=0$.
@@
Clearly, $f_{i}(\theta)\geq 0$ and hence $f(\theta)\geq 0$ for all $\theta$. Calculating the gradient and hessian for $f_{i}(\theta):=\exp(-y_{i}\langle x_{i},\theta\rangle)$,
$\nabla f_{i}(\theta)$ $=-\exp(-y_{i}\langle x_{i},\theta\rangle)y_{i}\,x_{i}$
@@
Hence, for all $i$, $f_{i}$ satisfies Assn. 5 with $L_{c}=0$ and $L_{g}=\max_{i}\left\|x_{i}\right\|$. Using Lemma 4, we conclude that $f(\theta)$ satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=8\,\max_{i}\left\|x_{i}\right\|_{2}^{2}$ and Assn. 3 with $\nu=8\,\max_{i}\left\|x_{i}\right\|$ and $\omega=0$. ∎
See 4
@@
From Mei et al. (2021, Lemma 2), we know that
$\left\|\nabla^{2}\ell(\theta)\right\|$ $=\left\|\nabla^{2}\langle\pi_{\theta},r\rangle\right\|\leq 3\,\left\|\nabla\langle\pi_{\theta},r\rangle\right\|=3\,\left\|\nabla\ell(\theta)\right\|$
Hence, the loss for the bandit problem satisfies Assn. 5 with $L_{c}=0$ and $L_{g}=3$. Using Lemma 4 with $n=1$, we can conclude the the loss for the bandit problem satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=72$ and Assn. 3 with $\nu=24$ and $\omega=0$. From Mei et al. (2020, Lemma 3), we know that,
@@
Hence, the loss for the bandit problem satisfies Assn. 4 with $\mu(\theta)=\pi_{\theta}(a^{*})$. ∎
@@
where $\pi^{*}$ is the optimal policy. $f(\theta)$ satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=8\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]^{2}\,S$,  Assn. 3 with $\nu=8\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\,\sqrt{S}$ and $\omega=0$ and  Assn. 4 with $\mu(\theta)=\frac{\min_{s}\pi_{\theta}(a^{*}(s)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}}\rho(s)}$ where $a^{*}(s)$ is the action that a deterministic optimal policy $\pi^{*}$ selects in state $s$.
@@
Assuming that the starting state distribution has full support, i.e. $\rho(s)>0$, from Mei et al. (2021, Lemma 6), we know that,